# DiCeption: A Generalist Diffusion Model for Visual Perceptual Tasks

### v2

@@ Abstract
Our primary goal here is to create a good, generalist perception model that can tackle multiple tasks, within limits on computational resources and training data. To achieve this, we resort to text-to-image diffusion models pre-trained on billions of images and introduce our visual generalist model: DiCeption. Our exhaustive evaluation metrics demonstrate that DiCeption effectively tackles multiple perception tasks, achieving performance on par with state-of-the-art models. We achieve results on par with SAM-vit-h using only 0.06% of their data (e.g., 600K vs. 1B pixel-level annotated images). Inspired by Wang et al. [109], DiCeption formulates the outputs of various perception tasks using color encoding; and we show that the strategy of assigning random colors to different instances is highly effective in both entity segmentation and semantic segmentation. Unifying various perception tasks as conditional image generation enables us to fully leverage pre-trained text-to-image models. Thus, DiCeption can be efficiently trained at a cost of orders of magnitude lower, compared to conventional models that were trained from scratch. When adapting our model to other tasks, it only requires fine-tuning on as few as 50 images and $\sim$1of its parameters. DiCeption provides valuable insights and a more promising solution for visual generalist models. Project webpage, and huggingface demo are available.
@@ 1 Introduction
Figure 1: With one single model, DiCeption solves multiple tasks without relying on any task-specific modules (rows 1 to 3). The red dots in the figure indicate the input points used for point-prompted segmentation. DiCeption preserves fine details in segmentation, such as hair (row 4). DiCeption supports both human pose estimation and semantic segmentation (row 5, 6). DiCeption can quickly adapt to new tasks by fine-tuning less than 1% of its parameters on as few as 50 images (row 7). For additional visualizations, please refer to Figures S1, S2, S4, S6, S7, S8, S9, S10, S11, S12, S13, S14 in the Appendix.
@@
Foundation models [49, 87, 118, 119, 117, 11, 7, 83, 75, 91, 6, 39], typically requiring extensive training on billions of data samples, play a pivotal role in their respective domains. In natural language processing (NLP), current foundation models [9, 99, 100, 26] have already demonstrated the potential to serve as versatile solutions, solving diverse fundamental tasks and with minimal fine-tuning needed for new tasks. This success can be attributed to the relatively small representational differences among various language tasks. However, in the domain of computer vision, task representations can differ substantially, and up to date, we still lack an effective approach to unify these distinct tasks. Consequently, existing vision foundation models usually excel at one single specific task, such as image segmentation [49, 87] or monocular depth estimation [118, 119, 117], because they are trained on data tailored exclusively to that task. Owing to the pronounced disparity in visual representations across tasks, coupled with the single-task specialization that characterizes current vision foundation models, fine-tuning these models for new tasks remains a formidable challenge. Although some efforts [12, 75, 39, 88] have been made to learn universal visual representations for more generalized vision foundation models, their performance still falls noticeably short compared to specialized models.
@@
Recent studies [109, 69, 68, 72, 2, 113] on visual generalist models are predominantly trained from scratch, often requiring substantial computational resources and large datasets to achieve good results. Unfortunately, the price of collecting a sufficiently large and high-quality multi-task dataset is substantial. Here, inspired by the success of diffusion models, we propose the hypothesis that leveraging their powerful priors can help mitigate the significant computational and data overhead for training powerful generalist models. While some existing works [47, 114, 38, 123, 93] have demonstrated that this is feasible in single-task scenarios, the potential of diffusion model priors in multi-task settings remains largely under-explored.
@@
In this paper, we successfully leverage the priors of diffusion models to achieve results on par or even better with the state-of-the-art models on various tasks with only minimal training data. We name our powerful visual generalist model DiCeption. For each task, we require substantially less data than specialized foundation models. For instance, compared to SAM segmentation, DiCeption  achieves comparable performance using a significantly smaller dataset of 600K samples, without any selective sampling. This is in sharp contrast to SAM which relies on a massive dataset of 1 billion pixel-level annotated samples. This demonstrates the efficiency and robustness of DiCeption, as it requires only 0.06% of the data used by SAM while still producing competitive results.
The concurrent work One Diffusion [53] proposes a unified model for image generation by treating different images as different views. However, their approach fails to distinguish between conditions and images, greatly limiting its effectiveness in perception tasks. More critically, their from-scratch training demands substantial data and computational resources. In contrast, our work focuses on image perception rather than generation, achieving on-par state-of-the-art performance in perception tasks. See more details in Appendix 2.4. DiCeption  highlights that the generative image priors lead to surprisingly more efficient and effective pathways to generalist image understanding models. The entire training process is remarkably stable, eliminating the need to design a complex training recipe. Even more notably, DiCeption is capable of quickly adapting to new tasks using as few as 50 training images and fine-tuning less than 1% of its parameters. Moreover, we show that previously ineffective strategies in generalist models such as task control via prompts [113] and random color assignments for segmentation masks [109] are viable in our framework. Additionally, we investigate the role of 1-step denoising in multi-task settings, a technique proven beneficial in single-task settings. DiCeption provides valuable insights for the design and training of strong visual generalist models.
@@
  • We introduce DiCeption, a generalist model capable of performing multiple visual perception tasks. Extensive experiments demonstrate that DiCeption can effectively address various tasks, achieving or par accuracy with state-of-the-art specialist models.
  • When trained on new tasks, DiCeption achieves high-quality results with minimal fine-tuning. Using only a small number of images (even with 50 images), the model can achieve exceptional performance by fine-tuning less than 1% of its parameters.
  • We conduct a thorough set of experiments exploring various aspects of multitask models, including network architectures, and 1-step diffusion. We believe that our experimental findings offer valuable insights and will significantly benefit future research.
@@ 2.1 Vision Foundation Models
Vision foundation models are models that are trained on large-scale datasets and demonstrate excellent performance within their trained domains. Vision foundation models now exist for a broad range of vision tasks, including monocular depth estimation [118, 119, 117, 7], object detection [11], segmentation [49, 87], multimodal tasks [83, 64], image and video generation [91, 28, 6], and more recently, emerging 3D models [105, 70]. While many works [110, 48, 58, 84, 133, 136] have sought to leverage the prior knowledge embedded in these models to tackle other tasks, such efforts often require complex network designs and intricate training strategies, typically transferring only to a limited number of tasks. Some foundation models [88, 39, 75, 12] emphasize representation learning, aiming to solve diverse downstream tasks by relying on generalized features. However, the results of these methods often fall short when compared with specialized foundation models. In contrast, our approach ensures consistent accuracy across multiple tasks while also enabling swift adaptation to new downstream tasks.
@@ 2.2 Diffusion Models
Diffusion models [28, 91, 6] have achieved remarkable success in image and video generation in recent years. The idea is to gradually add noise to the data and train a model to reverse this process, denoising step by step to generate the result. Recent diffusion models [28] utilize flow matching [63, 1, 66] and the DiT architecture [77], making them more scalable and efficient. Diffusion models have enabled a wide range of notable applications, including conditional image generation [130, 61, 124, 73, 82], image editing [8, 46, 116], story generation [107, 135], video generation [40, 35, 132, 122, 6, 50, 106], and video editing [13, 65, 14]. These successes underscore the substantial prior knowledge embedded in diffusion models.
Building on this insight, many studies [114, 38, 123, 47, 136] leverage diffusion models for downstream image understanding tasks. However, these approaches typically require separate fine-tuning for each individual task. Recently, we find several concurrent works [112, 53] also use diffusion models for multitask learning. Yet, these methods often involve complex network architectures and training procedures, and their evaluations tend to focus only on a very limited subset of image understanding results. In contrast, our DiCeption offers a simpler solution. We not only conduct detailed evaluations of our method across a variety of tasks but also demonstrate that the simplicity, paired with the inherent strengths of diffusion models, can be sufficient to deliver strong results without relying on overly complicated setups.
@@ 2.3 Multi-task Generalist Models
Recently, there has been a surge of interest in exploring visual multitask learning. Some approaches [109, 111] draw inspiration from in-context learning in NLP, adapting it for the visual domain. Others [69, 68, 72, 2] have advocated for sequence modeling methods, utilizing a transformer encoder-decoder architecture. In these approaches, different encoders map various tasks into a shared representation space, and distinct decoders are employed to transform tokens into the outputs specific to each task. However, these methods face notable limitations: they need to train a separate encoder and decoder for every individual task and they usually rely on substantial amounts of data to attain optimal performance.
The recent success of high-quality Vision Language Models (VLMs) [64] has also encouraged researchers to leverage them for building multitask models. Yet, these VLM-based methods [4, 104, 16, 67, 89, 59] typically focus on multimodal understanding tasks, such as image captioning, rather than general visual perception tasks. Meanwhile, some approaches [97, 132, 76] combine diffusion models with autoregressive models, focusing primarily on instruction-following image generation or editing tasks, rather than addressing image perception tasks. Although certain studies [52, 45, 17, 34] have tried to apply VLMs to more advanced semantic perception tasks, they struggle to establish a unified generalist visual model.
@@ 2.4 Compared with One Diffusion
The concurrent work, One Diffusion [53], addresses multi-task image generation, whereas our approach focuses on multi-task image understanding. We excel at performing a broader range of image understanding tasks with higher quality. While One Diffusion’s strategy of treating different images as different views benefits generation tasks, their failure to distinguish between conditions and images introduces harmful degrees of freedom for perception tasks, as illustrated in the red-highlighted regions of Figure S5. Specifically, when performing perception tasks, One Diffusion tends to generate an image similar to the original input, rather than the desired perceptual results.
@@ 3 Method
3.1 Preliminary
Recent diffusion models predominantly build upon flow matching methodologies [63, 1, 66] and the DiT architecture [77]. They aim to learn a velocity field that effectively maps samples from a source distribution to a target distribution. The training process involves minimizing the discrepancy between the model’s predicted velocity and the ground-truth velocity, which is typically formulated as:
${\rm Loss}={\mathbb{E}}_{\mathbf{z}_{0},t,c}\|v_{\theta}(\mathbf{z}_{t},t,c)-u(\mathbf{z}_{t})\|^{2}_{2},$ (1)
where $c$ is the condition, usually the text prompt. $\mathbf{z}_{0}$ is the clean image latent and $\mathbf{z}_{t}$ is the noisy one. The learned velocity field corresponds to an ordinary differential equation (ODE), such that during inference, a sample drawn from the source distribution can be transformed into the desired output by solving this ODE.
@@ 3.2 Unifying Task Representation into RGB Space
We focus on several of the most fundamental tasks in computer vision: monocular depth estimation, normal estimation, and segmentation. Segmentation, in particular, encompasses point-prompted segmentation, entity segmentation, and semantic segmentation. All these tasks can be unified within an RGB space, with the difference being the number of channels. For single-channel representations, such as depth maps and segmentation masks, we align them with RGB by repeating the channel data three times. For inherently three-channel representations, such as normal maps, we treat them directly as RGB images.
Entity segmentation is to segment every instance in an image but with no category. We assign each mask within an image a random color and merge them into a three-channel RGB mask. Painter found that assigning color randomly makes the model hard to optimize. However, we find this approach has no adverse impact on the training and enables the model to effectively learn to distinguish different instances by painting them with different colors. Each instance’s mask can be extracted from the RGB mask using clustering algorithms during post-processing without significant performance degradation. We also apply the random color assignment in semantic segmentation. Unlike traditional semantic segmentation, our method is capable of segmenting instances of the same semantic category. By default, We use KMeans for mask extraction.
@@ 3.3 DiCeption: A Unified Framework
Our model adopts the same architecture as SD3. We aim to keep the architecture as unchanged as possible, fully leveraging the pretrained prior knowledge. We use simple task prompts to direct the model to perform various tasks, such as “image2depth,” “image2normal,” and “image2segmentation.”
For point-prompted segmentation, a naive approach is directly painting points on the image. But this strategy is highly sensitive to the size of the points. If the painted points are too large, they can obscure small regions, causing segmentation to fail. Conversely, if the painted points are too small, the model may lose relevant point information after VAE downsampling and patchification. To address this, we introduce a minimal straightforward two-layer MLP $\Phi(\cdot)$ that enables the model to understand the point prompt.
Inspired by SAM, we apply sin-cos positional encoding to the point coordinates $p$, then pass them into the MLP $\Phi(\cdot)$ to produce point embeddings that match the dimension of the input hidden states. We use two learnable embeddings to indicate whether the embedding is a point embedding or not: $\xi_{p}$ for point embeddings and $\xi_{np}$ for non-point embeddings. The processed point embedding is summed with $\xi_{p}$. For other tasks, we simply use $\xi_{np}$ as the point embedding. During training, we randomly select 1–5 points to guide the segmentation. When the number of selected points is fewer than 5, we pad the point embeddings to a length of 5 with $\xi_{np}$. When performing tasks that do not require point input, the point embedding is simply a length-5 sequence, where each element is $\xi_{np}$. By denoting the final point embedding as $\xi$, this process is formulated as:
$\xi=\begin{cases}\operatorname{Concat}(\Phi(\operatorname{PE}(p))+\xi_{p},\xi_{np})&\text{if point segmentation}\\ \xi_{np}&\text{else}\end{cases}$ (2)
@@
DiCeption introduces two additional inputs based on SD3: the input image $\mathbf{x}^{\prime}$ and point embedding $\xi$. For the input image, we first apply VAE to down-sample it by a factor of 8, after which it is $2\times 2$ patchified into sequences. We denote this pre-processing as $\tau$. Subsequently, the task prompt token $\mathbf{e}$, point embedding $\xi$, noisy token $\mathbf{z}_{t}$, and input image token $\mathbf{z}^{\prime}$ are concatenated to form the complete input. We follow the flow matching loss in training SD3. During training, the loss is applied solely to the noisy tokens:
$\mathbf{z}_{0}$ $=\tau(\mathbf{x})$ (3)
$\mathbf{z}^{\prime}$ $=\tau(\mathbf{x}^{\prime})$

### v3

@@ Abstract
This paper’s primary objective is to develop a robust generalist perception model capable of addressing multiple tasks under constraints of computational resources and limited training data. We leverage text-to-image diffusion models pre-trained on billions of images and successfully introduce our DiCeption, a visual generalist model. Exhaustive evaluations demonstrate that DiCeption effectively tackles diverse perception tasks, even achieving performance comparable to SOTA single-task specialist models. Specifically, we achieve results on par with SAM-vit-h using only 0.06% of their data (e.g., 600K vs. 1B pixel-level annotated images). We designed comprehensive experiments on architectures and input paradigms, demonstrating that the key to successfully re-purposing a single diffusion model for multiple perception tasks lies in maximizing the preservation of the pre-trained model’s prior knowledge. Consequently, DiCeption can be trained with substantially lower computational costs than conventional models requiring training from scratch. Furthermore, adapting DiCeption to novel tasks is highly efficient, necessitating fine-tuning on as few as 50 images and approximately 1% of its parameters. Finally, we demonstrate that a subtle application of classifier-free guidance can improve the model’s performance on depth and normal estimation. We also show that pixel-aligned training, as is characteristic of perception tasks, significantly enhances the model’s ability to preserve fine details. DiCeption offers valuable insights and presents a promising direction for the development of advanced diffusion-based visual generalist models. Code and model: Homepage.
@@ 1 Introduction
Figure 1: With one single model, DiCeption solves 6 perception tasks without relying on any task-specific modules. The red dots in the figure indicate the input points used for interactive segmentation. DiCeption can quickly adapt to new tasks by fine-tuning less than 1% of its parameters on as few as 50 images. For additional visualizations, please refer to Figures S8, S11, S10, S15, S16, S17, S18, S19, S20, S21, S22 in the Appendix. We select Person as the instance segmentation example for the purpose of consistent visualization, which does not mean our method is limited to only human instances.
@@
Foundation models kirillov2023segment ; ravi2024sam ; yang2024depth ; yang2024depth2 ; yang2024depthvideo ; carion2020end ; bochkovskii2024depth ; radford2021learning ; oquab2023dinov2 ; rombach2022high ; blattmann2023stable ; he2022masked , typically requiring extensive training on billions of data samples, play a pivotal role in their respective domains. In natural language processing (NLP), current foundation models brown2020language ; touvron2023llama ; touvron2023llama2 ; dubey2024llama have already demonstrated the potential to serve as versatile solutions, solving diverse fundamental tasks and with minimal fine-tuning needed for new tasks. This success can be attributed to the relatively small representational differences among various language tasks. However, in the domain of computer vision, task representations can differ substantially, and up to date, we still lack an effective approach to unify these distinct tasks. Consequently, existing vision foundation models usually excel at one single specific task, such as image segmentation kirillov2023segment ; ravi2024sam or monocular depth estimation yang2024depth ; yang2024depth2 ; yang2024depthvideo , because they are trained on data tailored exclusively to that task. Owing to the pronounced disparity in visual representations across tasks, coupled with the single-task specialization that characterizes current vision foundation models, fine-tuning these models for new tasks remains a formidable challenge. Although some efforts caron2021emerging ; oquab2023dinov2 ; he2022masked ; ren2024dino have been made to learn universal visual representations for more generalized vision foundation models, their performance still falls noticeably short compared to specialized models.
@@
Recent studies wang2023images ; lu2022unified ; lu2024unified ; mizrahi20234m ; bachmann20244m ; xi2023dynamic on visual generalist models are predominantly trained from scratch, often requiring substantial computational resources and large datasets to achieve good results. Unfortunately, the price of collecting a sufficiently large and high-quality multi-task dataset is substantial. Here, inspired by the success of diffusion models, we propose the hypothesis that leveraging their powerful priors can help mitigate the significant computational and data overhead for training powerful generalist models. While some existing works ke2024repurposing ; xu2024diffusion ; he2024lotus ; ye2024stablenormal ; shao2024learning have demonstrated that this is feasible in single-task scenarios, the potential of diffusion model priors in multi-task settings remains largely under-explored.
@@
In this paper, we successfully leverage the priors of diffusion models to achieve results on par with the state-of-the-art models on various tasks with only minimal training data. We name our powerful visual generalist model DiCeption. For each task, we require substantially less data than specialized foundation models. For instance, compared to SAM segmentation trained on 1 billion pixel-level annotated samples, DiCeption achieves comparable performance using a significantly smaller dataset of 600K samples, without any training data cherry-picking.
More significantly, DiCeption highlights that the generative image priors lead to surprisingly more efficient and effective pathways to generalist image understanding models. We analyze a series of design choices for transferring one single modern diffusion model to multiple perception tasks, and identify that the key to successful transfer lies in preserving as much of the pretrained prior as possible, eliminating the need to design any complex module or training recipe. Even more notably, DiCeption is capable of quickly adapting to new tasks using as few as 50 training images and fine-tuning less than 1% of its parameters. We also demonstrate that pixel-level aligned training for perception tasks significantly enhances the model’s ability to preserve fine details and mitigates generated artifacts, which is of high significance for downstream applications. We believe DiCeption provides valuable insights for the design and training of strong diffusion-based visual generalist models.
@@
  • We introduce DiCeption, to the best of our knowledge, the first unified multi-task perception model with fully shared parameters that achieves quantitative performance comparable to specialized models while requiring significantly less data. E.g., we achieve competitive results with SAM-vit-h with only 0.06% of its data. We are capable of addressing six visual perception tasks within one single model.
  • This work offers a comprehensive experimental analysis elucidating the critical designs for effectively re-purposing diffusion models towards perception tasks, including architecture, input injection strategies and sampling timestep selection. Our findings establish that the preservation of the pretrained generative prior is paramount for achieving rapid adaptation and robust multi-task performance. Notably, DiT architectures are shown to be particularly conducive to this objective.
  • The proposed unified multi-task paradigm yields compelling advantages. For instance, DiCeption rapidly adapts to novel tasks in few-shot settings, demonstrating strong performance with as few as 50 training images and fine-tuning only 1% of parameters. Training on pixel alignment tasks significantly mitigates the artifacts often observed in other generative models for low-level image processing tasks such as image highlighting. Furthermore, the unified prediction space enables interactive segmentation to achieve matting-level accuracy.
@@ 2.1 Vision Foundation Models
Vision foundation models are models that are trained on large-scale datasets and demonstrate excellent performance within their trained domains. Vision foundation models now exist for a broad range of vision tasks, including monocular depth estimation yang2024depth ; yang2024depth2 ; yang2024depthvideo ; bochkovskii2024depth , object detection carion2020end , segmentation kirillov2023segment ; ravi2024sam , multimodal tasks radford2021learning ; liu2024visual , image and video generation rombach2022high ; esser2024scaling ; blattmann2023stable , and more recently, emerging 3D models wang2024dust3r ; ma2024you . While many works wang2024task ; khanna2024explora ; li2024matching ; rajivc2023segment ; zhong2024convolution ; zhu2024unleashing have sought to leverage the prior knowledge embedded in these models to tackle other tasks, such efforts often require complex network designs and intricate training strategies, typically transferring only to a limited number of tasks. Some foundation models ren2024dino ; he2022masked ; oquab2023dinov2 ; caron2021emerging emphasize representation learning, aiming to solve diverse downstream tasks by relying on generalized features. However, the results of these methods often fall short when compared with specialized foundation models. In contrast, our approach ensures consistent accuracy across multiple tasks while also enabling swift adaptation to new downstream tasks.
@@ 2.2 Diffusion Models
Diffusion models esser2024scaling ; rombach2022high ; blattmann2023stable have achieved remarkable success in image and video generation in recent years. The idea is to gradually add noise to the data and train a model to reverse this process, denoising step by step to generate the result. Recent diffusion models esser2024scaling utilize flow matching lipman2022flow ; albergo2022building ; liu2022flow and the DiT architecture peebles2023scalable , making them more scalable and efficient. Diffusion models have enabled a wide range of notable applications, including conditional image generation zhang2023adding ; li2024photomaker ; ye2023ip ; mou2024t2i ; qin2023unicontrol , image editing brooks2023instructpix2pix ; kawar2023imagic ; yang2023paint , story generation wang2024autostory ; zhou2024storydiffusion , video generation ho2022video ; guo2023animatediff ; zhao2024moviedreamer ; yang2024cogvideox ; blattmann2023stable ; kong2024hunyuanvideo ; wang2024framer , and video editing ceylan2023pix2video ; liu2024video ; chai2023stablevideo . These successes underscore the substantial prior knowledge embedded in diffusion models.
Building on this insight, many studies xu2024diffusion ; he2024lotus ; ye2024stablenormal ; ke2024repurposing ; zhu2024unleashing leverage diffusion models for downstream image understanding tasks. However, these approaches typically require separate fine-tuning for each individual task. Recently, we find several concurrent works wang2024lavin ; le2024diffusiongenerate also use diffusion models for multitask learning. Yet, these methods often involve complex network architectures and training procedures, and their evaluations tend to focus only on a very limited subset of image understanding results. In contrast, our DiCeption offers a simpler solution. We not only conduct detailed evaluations of our method across a variety of tasks but also demonstrate that the simplicity, paired with the inherent strengths of diffusion models, can be sufficient to deliver strong results without relying on overly complicated setups.
@@ 2.3 Multi-task Generalist Models
Recently, there has been a surge of interest in exploring visual multitask learning. Some approaches wang2023images ; wang2023seggpt draw inspiration from in-context learning in NLP, adapting it for the visual domain. Others lu2022unified ; lu2024unified ; mizrahi20234m ; bachmann20244m have advocated for sequence modeling methods, utilizing a transformer encoder-decoder architecture. In these approaches, different encoders map various tasks into a shared representation space, and distinct decoders are employed to transform tokens into the outputs specific to each task. However, these methods face notable limitations: they need to train a separate encoder and decoder for every individual task and they usually rely on substantial amounts of data to attain optimal performance.
The recent success of high-quality Vision Language Models (VLMs) liu2024visual has also encouraged researchers to leverage them for building multitask models. Yet, these VLM-based methods bai2023qwen ; wang2024qwen2 ; chen2024internvl ; lu2024deepseek ; ren2024pixellm ; li2025llamavid typically focus on multimodal understanding tasks, such as image captioning, rather than general visual perception tasks. Meanwhile, some approaches sun2024generative ; zhao2024moviedreamer ; pan2023kosmos combine diffusion models with autoregressive models, focusing primarily on instruction-following image generation or editing tasks, rather than addressing image perception tasks. Although certain studies lai2024lisa ; jiang2024chatrex ; cheng2024spatialrgpt ; guo2024regiongpt have tried to apply VLMs to more advanced semantic perception tasks, they struggle to establish a unified generalist visual model.
@@ 2.4 Compared with One Diffusion
The concurrent work, One Diffusion le2024diffusiongenerate , addresses multi-task image generation, whereas our approach focuses on multi-task image understanding. We excel at performing a broader range of image understanding tasks with higher quality. While One Diffusion’s strategy of treating different images as different views benefits generation tasks, their failure to distinguish between conditions and images introduces harmful degrees of freedom for perception tasks, as illustrated in the red-highlighted regions of Figure S14. Specifically, when performing perception tasks, One Diffusion tends to generate an image similar to the original input, rather than the desired perceptual results.
@@ 3 Method
3.1 Overview
Our methodology builds upon pre-trained text-to-image diffusion models esser2024scaling , steering perception tasks using text prompts. As shown in Figure 2, we concatenate the input image tokens, the noisy tokens, task prompt embeddings, and point embeddings for interactive segmentation along the token dimension. Training employs a flow matching loss esser2024scaling , exclusively computed on the noisy tokens. In inference, each denoising step refines only these noisy tokens, leaving all other conditioning tokens unchanged throughout the iterative denoising process.
Figure 2: We propose a generalist diffusion model solving multiple perception tasks, DiCeption. We select Person as the instance segmentation example for the purpose of consistent visualization, which does not mean our method is limited to only human instances. At each denoising step, the point embedding, input image latent, and task embedding remain fixed, while only the noise latent is updated.
@@ 3.2 Unifying Task Representation into RGB Space
We focus on several of the most fundamental tasks in computer vision: monocular depth estimation, normal estimation, human keypoint estimation and segmentation. Segmentation, in particular, encompasses interactive segmentation, entity segmentation, and instance segmentation. Our instance segmentation segments target instances with category name as input. All these tasks can be unified within an RGB space, with the difference being the number of channels. For single-channel representations, such as depth maps and segmentation masks, we align them with RGB by repeating the channel three times. For inherently three-channel representations, such as normal maps, we treat them directly as RGB images.
Entity segmentation is to segment every instance in an image but with no category. We assign each mask within an image a random color and merge them into a three-channel RGB mask. Painter wang2023images found that assigning color randomly makes the model hard to optimize. However, we find this approach has no adverse impact on the training and enables the model to effectively learn to distinguish different instances by painting them with different colors. Each instance’s mask can be extracted from the RGB mask using clustering algorithms during post-processing without significant performance degradation. We also apply the random color assignment in instance segmentation. Our method is capable of segmenting instances of the same semantic category. By default, we use KMeans for mask extraction.
@@ 3.3 DiCeption: A Unified Framework
Our model adopts the same architecture as SD3 esser2024scaling . We aim to keep the architecture as unchanged as possible, fully leveraging the pre-trained prior knowledge. To do so, we concatenate the input image tokens, noisy tokens, task embeddings, and point embeddings along the token dimension as input to the model. During training, the loss is computed only on the noisy tokens. Similarly, during inference, at each timestep, only the noisy tokens are updated, while the other tokens remain unchanged. We use simple task prompts to direct the model to perform various tasks, such as "image to depth", "image to normal", and "image to segmentation". An additional category name is provided in instance segmentation, such as "image to instance - cat".
Introduction of Point Embeddings
For point-prompted interactive segmentation, a naive approach is directly painting points on the image. But this strategy is highly sensitive to the size of the points. If the painted points are too large, they can obscure small regions, causing segmentation to fail. Conversely, if the painted points are too small, the model may lose relevant point information after VAE downsampling and patchification. To address this, we introduce a minimal straightforward two-layer MLP $\Phi(\cdot)$ that enables the model to understand the point prompt.
Inspired by SAM kirillov2023segment , we apply sin-cos positional encoding to the point coordinates $p$, then pass them into the MLP $\Phi(\cdot)$ to produce point embeddings that match the dimension of the input hidden states. We use two learnable embeddings to indicate whether the embedding is valid or not: $\xi_{p}$ for valid point embeddings and $\xi_{np}$ for invalid point embeddings. The processed point embedding is summed with $\xi_{p}$. For other tasks, we simply use $\xi_{np}$ as the point embedding. During training, we randomly select 1–5 points to guide the segmentation. When the number of selected points is fewer than 5, we pad the point embeddings to a length of 5 with $\xi_{np}$. When performing tasks that do not require point input, the point embedding is simply a length-5 sequence, where each element is $\xi_{np}$. By denoting the final point embedding as $\xi$, this process is formulated as:
[…]
@@
DiCeption introduces two additional inputs based on SD3: the input image $\mathbf{x}^{\prime}$ and point embedding $\xi$. For the input image, we first apply VAE to down-sample it by a factor of 8, after which it is $2\times 2$ patchified into sequences. We denote this pre-processing as $\tau$. Subsequently, the task prompt token $\mathbf{e}$, point embedding $\xi$, noisy token $\mathbf{z}_{t}$, and input image token $\mathbf{z}^{\prime}$ are concatenated along the token dimension to form the complete input. We follow the flow matching lipman2022flow ; albergo2022building ; liu2022flow loss in training SD3 esser2024scaling , which minimizes the discrepancy between the model’s predicted velocity $v$ and the ground-truth velocity $u$. During training, the loss is applied solely to the noisy tokens:
$\mathbf{z}_{0}=\tau(\mathbf{x})$ $,\mathbf{z}^{\prime}=\tau(\mathbf{x}^{\prime})$ (2)